OpenStudy (anonymous):

Solve. Show your work. A parallelogram has a base of 6.2 yd and a height of 7 yd. If the base and height are each multiplied by 4, what is the effect on the area of the parallelogram? Explain.

OpenStudy (phi):

Do you know the formula for the area of a parallelogram ?

OpenStudy (anonymous):

base times height

OpenStudy (phi):

A= b*h the base and height are each multiplied by 4, that means the new base is 4b (4 times bigger) the new height is 4h what is the new A ? (use the formula)

OpenStudy (anonymous):

173.6?

OpenStudy (anonymous):

694.4 I meant.

OpenStudy (phi):

I would leave it with the letters. just use 4b and 4h what do you get ?

OpenStudy (anonymous):

4b is 24.8 is 4h is 28

OpenStudy (phi):

yes. But there is an important idea we can get, if we leave the letters as letters what is the new area using 4b and 4h as the new base and new height?

OpenStudy (anonymous):

694.4?

OpenStudy (phi):

new A= 4*b * 4 * h when we multiply we can switch the order , so we can write is as new A = 4*4*bh or new A = 16 bh ok, so far?

OpenStudy (phi):

next, we notice this: the original area was A= bh (using the original "b" and "h") so new A= 16 bh can be written as new Area= 16 * (old Area) we don't need to use numbers for b and h, to see that whatever the numbers, the new area is 16 times bigger